OpenStudy (anonymous):

how do you solve this problem (x squared - 27x) = 90 degrees (right angle)

OpenStudy (anonymous):

\[x^2-27x=90\] \[(x-30)(x+3)=90\] etc

OpenStudy (anonymous):

ok well then does x=3 and x=-30 or is it just one answer?

OpenStudy (anonymous):

\[x ^{2} - 27x =90\] \[x ^{2} - 27x - 90 =0\] so (x-30)(x+3) = 0 so x = 30 OR x=-3 are correct answers

OpenStudy (anonymous):

THANKS SO MUCH!! I understand factoring now! :)

OpenStudy (anonymous):

THANKS SO MUCH!! I understand factoring now! :)

OpenStudy (anonymous):

THANKS SO MUCH!! I understand factoring now! :)